Each month, an American household generates an average of 28 pounds of newspaper for garbage and/or recycling. Assume this is approximately normally distributed and that the standard deviation is 2 pounds. If a household is selected at random, find the probability of it generating between 27 - 31 pounds per month.
Answer by Boreal(15235)

z=(x-mean)/sd
between z=(27=28)/2=-0.5 and z=(31-28)/2=+1.5
That probability is 0.6247